The sum of two times a number and negative thirteen is equal to the difference of the number and nine. Find the number. (the answer is 4 according to the answer key) I do not know how it got there..how is this problem written out?
Answer by rfer(16322)

2x-13=x-9
2x-x=-9+13
x=4